OpenStudy (anonymous):

Solve for x. -3x + 4 = -8

OpenStudy (anonymous):

Do you have a guess? If not I can tell you how to solve :)

OpenStudy (anonymous):

I am guessing it is -12

OpenStudy (anonymous):

Incorrect, I can tell you how to walk through.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The first step is to Subtract 4 from both sides. −3x+4−4=−8−4 −3x=−12

OpenStudy (gabbyalicorn):

−3x+4=−8 Subtract 4 from both sides. −3x+4−4=−8−4 −3x=−12 Divide both sides by -3. −3x/−3=−12/−3 x=4 Answer x=4

OpenStudy (anonymous):

Exactly what she said haha! Thanks! :D

OpenStudy (anonymous):

Thank you!! :D

OpenStudy (gabbyalicorn):

No problem! ^-^

OpenStudy (anonymous):

Answer x=4

OpenStudy (anonymous):

I wanna give both of you a medal

OpenStudy (anonymous):

If you didn't see it :P

OpenStudy (anonymous):

Thanks but give it to her

OpenStudy (anonymous):

ok